OpenStudy (anonymous):

I need help solving an=a1+(n-1)d and I'm suppose to solve for d. I need help please :)

OpenStudy (anonymous):

transpose a1 to left side.

OpenStudy (jdoe0001):

hmmm, that's just plain simplification, I assume you're not doing sequences evaluations

OpenStudy (phi):

The first thing to know is that "an" and "a1" are names and do not mean a times n or a times 1 an=a1+(n-1)d add -a1 to both sides an-a1 = a1-a1 + (n-1) d can you simplify the right side ?

OpenStudy (phi):

you should notice that a1 - a1 is 0 so you can simplify to an - a1 = (n-1) d now divide both sides by (n-1) can you do that ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Give it a try.

OpenStudy (anonymous):

got the answer thank u!

OpenStudy (phi):

to divide draw a big line underneath both sides then write (n-1) under the line on both sides